OpenStudy (experimentx):

Inverse Laplace transform of \( \cot^{-1}(s) \)

OpenStudy (experimentx):

using property no 30) http://tutorial.math.lamar.edu/Classes/DE/Laplace_Table.aspx let f(t) = sin(t), n=-1, \[ \sin(t) / t = -\int L(\sin(t)) ds = -\arctan(s)+c_1 = \cot^{-1}(s) + c_2\] they only differ by constant.